# SAFDNet: A Simple and Effective Network for Fully Sparse 3D Object Detection

### v1

@@ Abstract
LiDAR-based 3D object detection plays an essential role in autonomous driving. Existing high-performing 3D object detectors usually build dense feature maps in the backbone network and prediction head. However, the computational costs introduced by the dense feature maps grow quadratically as the perception range increases, making these models hard to scale up to long-range detection. Some recent works have attempted to construct fully sparse detectors to solve this issue; nevertheless, the resulting models either rely on a complex multi-stage pipeline or exhibit inferior performance. In this work, we propose SAFDNet, a straightforward yet highly effective architecture, tailored for fully sparse 3D object detection. In SAFDNet, an adaptive feature diffusion strategy is designed to address the center feature missing problem. We conducted extensive experiments on Waymo Open, nuScenes, and Argoverse2 datasets. SAFDNet performed slightly better than the previous SOTA on the first two datasets but much better on the last dataset, which features long-range detection, verifying the efficacy of SAFDNet in scenarios where long-range detection is required. Notably, on Argoverse2, SAFDNet surpassed the previous best hybrid detector HEDNet by 2.6% mAP while being 2.1$\times$ faster, and yielded 2.1% mAP gains over the previous best sparse detector FSDv2 while being 1.3$\times$ faster. The code will be available at https://github.com/zhanggang001/HEDNet.
@@ 3.2 Overall architecture
Since most high-performing hybrid detectors utilize the center-based head introduced by CenterPoint [4], we adopt a similar head for SAFDNet. However, as the designs for classification and regression in CenterPoint are tailored for dense feature maps, we have made some adjustments to accommodate sparse features. For more details, please refer to Section 3.4.
@@ 3.3 Adaptive feature diffusion
SWFormer [${}^{\dagger}$/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
VoxelNeXt [${}^{\dagger}$ 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
FSDv1 [${}^{\dagger}$ 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
FSDv2 [${}^{\dagger}$ 81.8/79.5 75.6/73.5 79.8/79.3 71.4/71.0 84.8/79.7 77.4/72.5 80.7/79.6 77.9/76.8
SAFDNet (${}^{\dagger}$ 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
@@
FSDv1 [${}^{\dagger}$ 80.4/78.2 74.4/72.4 82.7/82.3 74.4/74.1 82.9/77.9 75.9/71.3 75.6/74.4 72.9/71.8
FSDv2 [${}^{\dagger}$ 81.1/79.0 75.4/73.3 82.4/82.0 74.4/74.0 83.8/78.9 77.4/72.8 77.1/76.0 74.3/73.2
SAFDNet (${}^{\dagger}$ 81.9/79.8 76.5/74.6 83.9/83.5 76.6/76.2 84.3/79.8 78.4/74.1 77.5/76.3 74.6/73.4
Table 2: Comparison with prior methods on the Waymo Open dataset. Metrics: mAP/mAPH (%$\uparrow$ for the overall results, and AP/APH (%$\uparrow$ for each category. ${}^{\dagger}$ represents a fully sparse detector, the same as below. All models were trained under single-frame setting.
@@
VoxelNeXt [${}^{\dagger}$ 68.7 63.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
FSDv2 [${}^{\dagger}$ 70.4 64.7 84.4 57.3 75.9 44.1 28.5 86.9 69.5 57.4 72.9 73.6
SAFDNet (${}^{\dagger}$ 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
@@
VoxelNeXt [${}^{\dagger}$ 70.0 64.5 84.6 53.0 64.7 55.8 28.7 85.8 73.2 45.7 79.0 74.6
FSDv2 [${}^{\dagger}$ 71.7 66.2 83.7 51.6 66.4 59.1 32.5 87.1 71.4 51.7 80.3 78.7
SAFDNet (${}^{\dagger}$ 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
@@ 3.4 Sparse detection head
For the detection head, we adhere to most design principles of CenterPoint but make some adjustments to accommodate sparse features. Despite employing adaptive feature diffusion, covering all object centers remains challenging, particularly for extremely large objects. For classification training, we calculate a Gaussian heatmap for each object using the distance from each voxel center to the object center. Different from CenterPoint, we normalize the heatmap values with the maximum value for each object to avoid gradient vanishing. For regression, we compute the loss based on predictions from voxels features nearest to object centers.
@@ 4.3 Comparison with state-of-the-art methods
VoxelNeXt [${}^{\dagger}$ 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
FSDv1 [${}^{\dagger}$ 28.2 68.1 40.9 59.0 29.0 38.5 41.8 42.6 39.7 26.2 49.0 38.6 20.4 30.5 14.8 41.2 26.9 11.9 5.9 13.8 33.4 21.1 0.0 9.5 7.1 14.0 9.2
FSDv2 [${}^{\dagger}$ 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
SAFDNet (${}^{\dagger}$ 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
@@
VoxelNeXt [${}^{\dagger}$ 70.1 15.7 0.9$\times$ 30.7 19.6 2.7$\times$ 6.2G
FSDv2 [${}^{\dagger}$ 73.5 10.3 0.6$\times$ 37.6 11.5 1.6$\times$ 8.6G
${}^{\dagger}$(Ours) 73.9 20.2 1.2$\times$ 39.7 15.1 2.1$\times$ 7.3G
@@
(a) A step-by-step ablation.
(b) Different types of feature diffusion.
@@
$\alpha$ Waymo Open Argoverse2 mAPH FLOPs mAP FLOPs 0.0 71.5 90G 36.4 34G 0.5 72.7 100G 37.3 38G 1.0 73.3 108G 37.8 45G 2.0 73.0 153G 37.8 67G
(a) A step-by-step ablation.
@@ Appendix A Implementations details
We trained all models with a batch size of 16 on 8 RTX 4090 GPUs. We present more details as follows.
@@
We set the voxel size to (0.08m, 0.08m, 0.15m), and the detection range to [-75.52m, 75.52m] in X and Y axes, and [-2m, 4m] in Z axis. We trained SAFDNet for 24 epochs on the training dataset and reported the evaluation results on the validation set to compare with previous methods. For results on the test set, we trained the model on both the training and validation sets. We employed the Adam [46] optimizer with a one-cycle learning rate policy, and set the weight-decay to 0.05, and the max learning rate to 0.003. We adopted the faded training strategy in the last epoch. During inference, we applied class-specific NMS with an IoU threshold of 0.75, 0.6 and 0.55 for vehicle, pedestrian, and cyclist, respectively. There are three category groups: group 1 includes vehicle; group 2 includes pedestrian and cyclist; group 3 is the background. The kernel size K for feature diffusion for the three groups are set to 7, 3, and 3, respectively.
@@
We set the voxel size to (0.075m, 0.075m, 0.2m), and the detection range to [-54m, 54m] in X and Y axes, and [-5m, 3m] in Z axis. We trained SAFDNet for 20 epochs on both training and validation sets and reported the evaluation results on the test set to compare with previous methods. We employed the Adam optimizer with a one-cycle learning rate policy, and set the weight-decay to 0.1, the momentum to [0.85, 0.95], and the max learning rate to 0.001. The faded strategy was used during the last 5 epochs. We set the query number of detection head to 300 and did not use any test-time augmentation. There are four category groups: group 1 includes bus and trailer; group 2 includes car, truck, and construction vehicle; group 3 includes motorcycle, bicycle, traffic cone, pedestrian, and barrier; group 4 is the background. The kernel size K for feature diffusion for the four groups are set to 15, 9, 3, and 3, respectively.
@@
We set the voxel size to (0.1m, 0.1m, 0.2m), and the detection range to [-200m, 200m] in X and Y axes, and [-4m, 4m] in Z axis. We trained SAFDNet for 24 epochs on the training set and reported the results on the validation set to compare with other methods. We employed the Adam optimizer with a one-cycle learning rate policy, and set the weight-decay to 0.05, and the max learning rate to 0.003. We adopted the faded training strategy in the last epoch. There are four category groups: group 1 includes large vehicle, bus, box truck, truck, truck cab, vehicular trailer, school bus, articulated bus, and message board trailer; group 2 includes regular vehicle; group 3 includes the other object categories; group 4 is the background. The kernel size K for feature diffusion for the four groups are set to 13, 7, 3, and 3, respectively.

### v3

@@ Abstract
LiDAR-based 3D object detection plays an essential role in autonomous driving. Existing high-performing 3D object detectors usually build dense feature maps in the backbone network and prediction head. However, the computational costs introduced by the dense feature maps grow quadratically as the perception range increases, making these models hard to scale up to long-range detection. Some recent works have attempted to construct fully sparse detectors to solve this issue; nevertheless, the resulting models either rely on a complex multi-stage pipeline or exhibit inferior performance. In this work, we propose a fully sparse adaptive feature diffusion network (SAFDNet) for LiDAR-based 3D object detection. In SAFDNet, an adaptive feature diffusion strategy is designed to address the center feature missing problem. We conducted extensive experiments on Waymo Open, nuScenes, and Argoverse2 datasets. SAFDNet performed slightly better than the previous SOTA on the first two datasets but much better on the last dataset, which features long-range detection, verifying the efficacy of SAFDNet in scenarios where long-range detection is required. Notably, on Argoverse2, SAFDNet surpassed the previous best hybrid detector HEDNet by 2.6% mAP while being 2.1$\times$ faster, and yielded 2.1% mAP gains over the previous best sparse detector FSDv2 while being 1.3$\times$ faster. The code will be available at https://github.com/zhanggang001/HEDNet.
@@ 3.2 Overall architecture
Since most high-performing hybrid detectors utilize the center-based head introduced by CenterPoint [4] or TransFusion-L [5], we adopt a similar head in SAFDNet. However, as the designs of these detection heads are tailored for dense feature maps, we have made some adjustments to accommodate sparse features. For more details, please refer to Section 3.4.
@@ 3.3 Adaptive feature diffusion
SWFormer [11] -/- -/- 77.8/77.3 69.2/68.8 80.9/72.7 72.5/64.9 -/- -/-
VoxelNeXt [13] 78.6/76.3 72.2/70.1 78.2/77.7 69.9/69.4 81.5/76.3 73.5/68.6 76.1/74.9 73.3/72.2
FSDv1 [10] 79.6/77.4 72.9/70.8 79.2/78.8 70.5/70.1 82.6/77.3 73.9/69.1 77.1/76.0 74.4/73.3
FSDv2 [12] 81.8/79.5 75.6/73.5 79.8/79.3 71.4/71.0 84.8/79.7 77.4/72.5 80.7/79.6 77.9/76.8
SAFDNet (ours) 81.8/79.9 75.7/73.9 80.6/80.1 72.7/72.3 84.7/80.4 77.3/73.1 80.0/79.0 77.2.76.2
@@
FSDv1 [10] 80.4/78.2 74.4/72.4 82.7/82.3 74.4/74.1 82.9/77.9 75.9/71.3 75.6/74.4 72.9/71.8
FSDv2 [12] 81.1/79.0 75.4/73.3 82.4/82.0 74.4/74.0 83.8/78.9 77.4/72.8 77.1/76.0 74.3/73.2
SAFDNet (ours) 81.9/79.8 76.5/74.6 83.9/83.5 76.6/76.2 84.3/79.8 78.4/74.1 77.5/76.3 74.6/73.4
Table 2: Comparison with prior methods on the Waymo Open dataset. Metrics: mAP/mAPH (%$\uparrow$ for the overall results, and AP/APH (%$\uparrow$ for each category. represents a fully sparse detector, the same as below. All models were trained under single-frame setting.
@@
VoxelNeXt [13] 68.7 63.5 83.9 55.5 70.5 38.1 21.1 84.6 62.8 50.0 69.4 69.4
FSDv2 [12] 70.4 64.7 84.4 57.3 75.9 44.1 28.5 86.9 69.5 57.4 72.9 73.6
SAFDNet (Ours) 71.0 66.3 87.6 60.8 78.0 43.5 26.6 87.8 75.5 58.0 75.0 69.7
@@
VoxelNeXt [13] 70.0 64.5 84.6 53.0 64.7 55.8 28.7 85.8 73.2 45.7 79.0 74.6
FSDv2 [12] 71.7 66.2 83.7 51.6 66.4 59.1 32.5 87.1 71.4 51.7 80.3 78.7
SAFDNet (Ours) 72.3 68.3 87.3 57.3 68.0 63.7 37.3 89.0 71.1 44.8 84.9 79.5
@@ 3.4 Sparse detection head
We adhere to most design principles of CenterPoint (Waymo Open and Argoverse2) and TransFusion-L (nuScenes), but make some adjustments to accommodate sparse features. Despite employing adaptive feature diffusion, covering all object centers remains challenging, particularly for extremely large objects. Therefore, we adopt the diffused voxel features nearest to object centers as object representations. For classification training, we calculate a Gaussian heatmap for each object based on the distance from each voxel center to the object center. We normalize the heatmap values by the maximum value for each object to avoid gradient vanishing.
@@ 4.3 Comparison with state-of-the-art methods
VoxelNeXt [13] 30.7 72.7 38.8 63.2 40.2 40.1 53.9 64.9 44.7 39.4 42.4 40.6 20.1 25.2 19.9 44.9 20.9 14.9 6.8 15.7 32.4 16.9 0.0 14.4 0.1 17.4 6.6
FSDv1 [10] 28.2 68.1 40.9 59.0 29.0 38.5 41.8 42.6 39.7 26.2 49.0 38.6 20.4 30.5 14.8 41.2 26.9 11.9 5.9 13.8 33.4 21.1 0.0 9.5 7.1 14.0 9.2
FSDv2 [12] 37.6 77.0 47.6 70.5 43.6 41.5 53.9 58.5 56.8 39.0 60.7 49.4 28.4 41.9 30.2 44.9 33.4 16.6 7.3 32.5 45.9 24.0 1.0 12.6 17.1 26.3 17.2
SAFDNet (Ours) 39.7 78.5 49.4 70.7 51.5 44.7 65.7 72.3 54.3 49.7 60.8 50.0 31.3 44.9 24.7 55.4 31.4 22.1 7.1 31.1 42.7 23.6 0.0 26.1 1.4 30.2 11.5
@@
VoxelNeXt [13] 70.1 15.7 0.9$\times$ 30.7 19.6 2.7$\times$ 6.2G
FSDv2 [12] 73.5 10.3 0.6$\times$ 37.6 11.5 1.6$\times$ 8.6G
SAFDNet(Ours) 73.9 20.2 1.2$\times$ 39.7 15.1 2.1$\times$ 7.3G
@@
$\alpha$ Waymo Open Argoverse2
mAPH FLOPs mAP FLOPs
0.0 71.5 90G 36.4 34G
0.5 72.7 100G 37.3 38G
1.0 73.3 108G 37.8 45G
2.0 73.0 153G 37.8 67G
@@ Appendix A Implementations details
We trained all models with a batch size of 16, employed the Adam [46] optimizer with a one-cycle learning rate policy, and set the weight decay to 0.05 and the maximum learning rate to 0.003. More details are presented below.
@@
We adopt a detection head derived from CenterPoint. We set the voxel size to (0.08m, 0.08m, 0.15m), and the detection range to [-75.52m, 75.52m] in X and Y axes, and [-2m, 4m] in Z axis. We trained SAFDNet for 24 epochs on the training dataset and reported the evaluation results on the validation set to compare with previous methods. For results on the test set, we trained the model on both the training and validation sets. We adopted the faded training strategy in the last epoch. During inference, we applied class-specific NMS with an IoU threshold of 0.75, 0.6 and 0.55 for vehicle, pedestrian, and cyclist, respectively. There are three category groups: group 1 includes vehicle; group 2 includes pedestrian and cyclist; group 3 is the background. The kernel size K for feature diffusion for the three groups are set to 7, 3, and 3, respectively.
@@
We adopt a detection head derived from TransFusion-L.We set the voxel size to (0.3m, 0.3m, 8.0m), and the detection range to [-54m, 54m] in X and Y axes, and [-5m, 3m] in Z axis. We trained SAFDNet for 20 epochs on both training and validation sets and reported the evaluation results on the test set to compare with previous methods. The faded strategy was used during the last 4 epochs. We set the query number of detection head to 300 (test only) and did not use any test-time augmentation. There are four category groups: group 1 includes bus and trailer; group 2 includes car, truck, and construction vehicle; group 3 includes motorcycle, bicycle, traffic cone, pedestrian, and barrier; group 4 is the background. The kernel size K for feature diffusion for the four groups are set to 15, 9, 7, and 7, respectively.
@@
We adopt a detection head derived from CenterPoint. We set the voxel size to (0.1m, 0.1m, 0.2m), and the detection range to [-200m, 200m] in X and Y axes, and [-4m, 4m] in Z axis. We trained SAFDNet for 24 epochs on the training set and reported the results on the validation set to compare with other methods. We adopted the faded training strategy in the last epoch. There are four category groups: group 1 includes large vehicle, bus, box truck, truck, truck cab, vehicular trailer, school bus, articulated bus, and message board trailer; group 2 includes regular vehicle; group 3 includes the other object categories; group 4 is the background. The kernel size K for feature diffusion for the four groups are set to 13, 7, 3, and 3, respectively.